which of the following could be the equation of a line which is passing through x-2y+3=0 and 2x-2y+6=0 and having a slope of -3 ?
a) y=-3x+3
b) y=-3x-6
c) y=-3
d) y=-3x-9
e) y=-3x+6

Answer by Alan3354(69443)

which of the following could be the equation of a line which is passing through x-2y+3=0 and 2x-2y+6=0 and having a slope of -3 ?
a) y=-3x+3
b) y=-3x-6
c) y=-3
d) y=-3x-9
e) y=-3x+6
===================
All 5 of the lines pass thru x-2y+3=0 and 2x-2y+6=0.
All except c have a slope of -3
--------------------
If you mean which one passes thru the intersection of x-2y+3=0 and 2x-2y+6=0, it's d) y = -3x-9